Find the thirtieth term of the following sequence.
-6, -4, -2, 0, ...
A.) 52
B.) 64
C.) 66

Answer by LinnW(1048)

a formula for this sequence is
2s -8 where s represents the term of the sequence so
for example, the first element is 2(1) - 8 = -6 ,
2(2) - 8 = -4
So the 13th term is 2(13) - 8 = 26 - 8 = 18